14 points SerPSET9 32.P034. A 10.5-V battery, a 4.93-? resistor, and a 9.3-H inductor are connected in series. After the current in the circuit has reached its maximum value, calculate the following (a) the power being supplied by the battery (b) the power being delivered to the resistor (c) the power being delivered to the inductor d) the energy stored in the magnetic field of the inductor Need Help?Read It
Answer #1

Given , V = 10.5 V ,R = 4.93 ohms , L = 9.3 H

Maximum current is at :
i = V/R = 10.5/4.93 = 2.13A
a) P(supplied)= V*i = 2.13*10.5 = 22.365 W.
b) P(resistor) = i2 * R = 2.132* 4.93 = 22.367 W.
c) P(inductor) = 0
d ) W= L*i2/ 2 = 9.3* 2.132/ 2 = = 21.1 J